40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jun 2 18:09:32 2010, steve, Update, PEM, seismometers off of linoleum floor 
    Reply  Wed Jun 2 21:21:43 2010, steve, Update, PEM, seismometers off of linoleum floor 
       Reply  Thu Jun 3 22:17:05 2010, rana, Update, PEM, seismometers off of linoleum floor Untitled.pnga.pngUntitled.png
          Reply  Fri Jun 4 23:52:57 2010, rana, Update, PEM, seismometers off of linoleum floor Untitled.pnghuddlez.png
Message ID: 3037     Entry time: Wed Jun 2 18:09:32 2010     Reply to this: 3039
Author: steve 
Type: Update 
Category: PEM 
Subject: seismometers off of linoleum floor 

Steve for Nancy,

Seismometer interface box ac power was turned off, Guralps disconnected and moved. Ranger locked, moved and released. Nancy will describe the rest soon.

 

ELOG V3.1.3-